OpenStudy (anonymous):

The Pythagorean Theorem shows the relationship between the sides of a right triangle and the hypotenuse. The formula for the Pythagorean Theorem is a2 + b2 = c2, where a and b are the sides and c is the hypotenuse. A. Write a polynomial expression for c2. (

OpenStudy (anonymous):

There's probably a picture.

OpenStudy (anonymous):

sry guuurl. Requires login.

OpenStudy (anonymous):

can you see it ?

OpenStudy (ajprincess):

(x+3)^2+(x+2)^2=c^2 can u simplify?

OpenStudy (anonymous):

im not sure ,

OpenStudy (anonymous):

Give it a try, and ask any questions that come up.

OpenStudy (anonymous):

okay ,

OpenStudy (anonymous):

i dont get it .

OpenStudy (anonymous):

"I don't get it" isn't a question.

OpenStudy (anonymous):

how do you simplifiy it from (x+3)^2 +(x+2)^2 = c^2 ?

OpenStudy (ajprincess):

using this form (a+b)^2=a^2+b^2+2ab expand (x+3)^2 and (x+2)^2. Then add the like terms to simplify. see if u can do nw.

OpenStudy (anonymous):

im sorry , but i just cant understand this.

OpenStudy (ajprincess):

\[(a+b)^2=a^2+2ab+b^2\] in \[(x+3)^2\] u hav a=x and b=3 applying this in \[a^2+2ab+b^2\] u get \[x^2+2*x*3+3^2\] \[=x^2+6x+9\] similarly can u do for \[(x+2)^2\]?

OpenStudy (anonymous):

yes.

OpenStudy (ajprincess):

vat do u get for that dr?

OpenStudy (anonymous):

idk .